A father is now only three times as old as his son but eight years ago the father was five times as old as his son. Find their present ages.
Answer by ankor@dixie-net.com(22740)

Let f = father's present age
let s = son's present age
:
A father is now only three times as old as his son
f = 3s
:
eight years ago the father was five times as old as his son.
f - 8 = 5(s-8)
f - 8 = 5s - 40
f = 5s - 40 + 8
f = 5s - 32
Replace f with 3s from the 1st statement
3s = 5s - 32
32 = 5s - 3s
32 = 2s
s = 16 is son's present age
I'll let you find father's age, check your solutions in the 2nd statement



Find their present ages.